how many grams of oxygen gas must be reacted with excess sulfur dioxide to produce 87,1 g of SO3?


1
Expert's answer
2022-04-14T11:26:04-0400

2SO2 + O2 = 2SO3

According to the equation, n (O2) = 1/2 x n (SO3)

n = m / M

M (SO2) = 64.1 g/ mol

M (O2) = 32 g/ mol

M (SO3) = 80.1 g/ mol

n (SO3) = m / M = 87.1 / 80.1 = 1.1 mol

n (O2) = 1/2 x n (SO3) = 1/2 x 1.1 = 0.54 mol

m (O2) = n x M = 0.54 x 32 = 17.4 g